OpenStudy (anonymous):

-x+3y+λ(x+2y+4)=0 p(-1,1)

OpenStudy (anonymous):

um what does the p-11 thing mean?

OpenStudy (anonymous):

what method do u want me to use to solve the equation can u specify?

OpenStudy (anonymous):

uhm i dont know to say it in english but ill try it says The straight line which passes through the point (-1,1) is A) 2x-y-5 = 0, b)-9x +7 y-16 = 0 c) 8x +11 +3 y = 0 d) x +4 y-​​1 = 0